OpenStudy (anonymous):

PLEASE HELP IM DESPERATE!!!!! Solve the following system of equations. x + y = 3 2x + 2y = 6

OpenStudy (anonymous):

Let's use the substitution method.

OpenStudy (anonymous):

{x + y = 3 {2x + 2y = 6 Let's solve for the x from the first equation. x = 3 - y Now we substitute this into the other equation: 2(3 - y) + 2y = 6 Can you try to solve for y? [sorry for the late reply]

OpenStudy (anonymous):

Once you solve for y you input the y-value into any of the equations and solve for x. Then, once you have the x-value, you put them together [(x, y)] and you have your solution. :)

OpenStudy (anonymous):

Thank you!!!!